OpenStudy (anonymous):

graph the equation by using the slope and y intercept y=1/8x+5 I have to graph this i can graph (0,5) but i have no clue how to graph 1/8 ??

OpenStudy (ybarrap):

Plug in various values for x and see what y pops out, then you have several (x,y) coordinates that you can plot. You really only need two points to connect the dots and you're done

OpenStudy (anonymous):

you have the y-intercept; if you want the x-intercept then simply let y=0 and solve for x:\[0=\frac{1}{8}x+5\]\[-5=\frac{1}{8}x\]\[x=-40\]then use the point (-40,0) along with the previously found (0,5). you will have to use correct relative position graphing which is expected in algebra classes.

OpenStudy (anonymous):

if you don't need the x-intercept, just choose any convenient x-value.. any mutliple of 8 would be easy arithmetically withe the 1/8

OpenStudy (anonymous):

ok so i take the multiple of 8, and i do not remember how to solve my notes are crappy

OpenStudy (anonymous):

for example. let x=8, then y=(1/8)(8)+5=1+5=6 => (8,6) or let x=16, then y=(1/8)(16)+5=2+5=7 => (16,7) etc.

OpenStudy (anonymous):

so if i use 16, then my fraction would b 16/8-2+5= (16,3) ????

OpenStudy (anonymous):

where did the "-2" come from? it would be 16/8 + 5 = 2+5 = 7 (see my previous post)

OpenStudy (anonymous):

i divided 16/8 for 2 lol???

OpenStudy (anonymous):

OOOHHHHH i see now

OpenStudy (anonymous):

its always easy when you see it

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

i have one more question